4 = 2/3x + 9+ 1/3x
Solve the equation by simplifiing both sides and isolating the variable

Found 2 solutions by edjones, stanbon:
Answer by edjones(8007)

4 = 2x/3 + 9+ x/3
3*4=3*2x/3+3*9+3*x/3
12=2x+27+x no fractions!
3x+27=12
3x+27-27=12-27
3x= -15
3x/3= -15/3
x= -5
Ed

Answer by stanbon(75887)

4 = 2/3x + 9+ 1/3x
4 = x + 9
x = -5
-------------
Cheers,
Stan H.